wer: hours
In a data distribution, the first quartile, the median and the means are 30.8, 48.5 and 42.0 respectively. If the coefficient skewness is −0.38 ,
1
Expert's answer
2020-06-08T19:15:37-0400

approximate value of the third quartile:

z value for Q1(25%)=-0.67

-0.67=(Q1-μ)/σ=-11.2

σ=-11.2/-0.67=16.7

3*IQR = 4*σ

IQR =4*σ/3=22.2

IQR =Q3-Q1, so Q3=IQR+Q1=22.2+30.8=53.0

approximate value of the variance

VAR=σ2=279